OpenStudy (anonymous):

okay.. simplify this.. r/r-1 - r²/r² .... good luck lol

OpenStudy (dumbcow):

well, r^2/r^2 = 1

OpenStudy (anonymous):

r/r-1 - 1 r-r-1/r-1 -1/r-1

OpenStudy (anonymous):

r/r-1 - r²/r² r²/r²=1 (r/r-1)-1 (r/r-1)-(r-1/r-1) r-r-1/r-1 -1/r-1

OpenStudy (anonymous):

where did the second r come from din?

OpenStudy (dumbcow):

\[\frac{r}{r-1} - 1 = \frac{r}{r-1} -\frac{r-1}{r-1} = \frac{r-(r-1)}{r-1} = \frac{1}{r-1}\]

OpenStudy (anonymous):

I'm sorry, how are you getting r-1/r-1?

OpenStudy (dumbcow):

just adding it, but because it equals 1 it doesn't change expression's value

OpenStudy (anonymous):

ooo

OpenStudy (dumbcow):

thats the only way to combine fractions, get a common denominator

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

thanks for your explanations dumbcow.

OpenStudy (dumbcow):

no problem, thats how you learn thanks for wanting them :)